OpenStudy (anonymous):

If X = 5 and Y = 3, what does Z equal? 2X + 2Z = 10Y 10 20 40 3 1/3 @satellite73 i need help

OpenStudy (anonymous):

\[2x+2a=10y\]?

OpenStudy (anonymous):

\[2\times 5+2\times Z=10\times 3\] if i am reading this correctly

OpenStudy (anonymous):

that means \[10+2Z=30\] solve for Z this is way easier then the previous questions, so maybe i am reading it wrong

OpenStudy (anonymous):

\[2Z=20\\Z=10\]is my answer

OpenStudy (anonymous):

oh yeahh lol i got the same thing i thought it was gonna be way harder than this

OpenStudy (anonymous):

how did we go from translating functions to this easy algebra?

OpenStudy (anonymous):

i just started taking this course lol so the beginning is very easy and throughout the whole thing it will get harder than ever :P

OpenStudy (anonymous):

What would be the value of angle a in the diagram found below of two parallel lines? 40 degrees 90 degrees 60 degrees 120 degrees cannot be determined

OpenStudy (anonymous):

would i be 40 degrees as well? @satellite73

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the internal angles add up to 180, you have a right angle at 90 and another angle of 40 leaving \[180-90-40=5\]

OpenStudy (anonymous):

ooops \[180-90-40=50\]

OpenStudy (anonymous):

its not in the answer choices.

OpenStudy (anonymous):

oooh i should pay attention it want the angle A, not the other angle of the triangle you are right, it is also 40 sorry

OpenStudy (anonymous):

lol okayy! thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

What is the value of the radius of the following circle with an area of 154 cm^2? 4.9 cm 7.0 cm 25 cm 68 cm

OpenStudy (anonymous):

\[A=\pi r^2\] \[154=\pi r^2\\ \frac{154}{\pi}=r^2\\ \sqrt{\frac{154}{\pi}}=r\]

OpenStudy (anonymous):

the answer is A?

OpenStudy (anonymous):

i get 7 http://www.wolframalpha.com/input/?i=sqrt%28154%2Fpi%29

OpenStudy (anonymous):

no i think it is B, 7

OpenStudy (anonymous):

oh yeahh hahah i forgot to square root oops lol :P

OpenStudy (anonymous):

Find the mean and median of the following data set.98, 87, 92, 79, 65, 91, 80, 92, 85, 86. Mean: 76.9, Median: 65 Mean: 84.5, Median: 78 Mean: 85.5, Median: 86.5 Mean: 85.5, Median: 86

OpenStudy (anonymous):

wow talk about donkey work for wolfram

OpenStudy (anonymous):

looks like mean is \(85.5\) now lets find the median

OpenStudy (anonymous):

looks like C

OpenStudy (anonymous):

Sanjay has quiz scores of 96%, 89%, 92%, 77%, 100% and 87%. What score does he need to make on the final quiz to have an overall average of 91%? 96% 97% 98% 99%

OpenStudy (anonymous):

i got A @satellite73

OpenStudy (anonymous):

\[\frac{96+89+92+77+100+87+x}{7}=91\]

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i= \frac{96%2B89%2B92%2B77%2B100%2B87%2Bx}{7}%3D91 looks like 96

OpenStudy (anonymous):

hmmm link may not work i get 96 so A